OpenStudy (kitkat1):

What is the distance from (–4, 0) to (2, 5)? Round your answer to the nearest hundredth. 8.17 8.58 10.4 7.81

OpenStudy (campbell_st):

just use pythaoras' theorem  so you have a right triangle with sides 5 and 6 an you need to find the hypotenuse

OpenStudy (kitkat1):

i ended up with 7.81 answer

OpenStudy (kitkat1):

Can u help me with one more?

OpenStudy (campbell_st):

just post your question in the normal way, there are lots of people who can help

OpenStudy (wmj259):

Distance=\[\sqrt{\left( x _{2}-x _{1} \right)^{2}+\left( y _{2}-y _{1} \right)^{2}}\]

OpenStudy (wmj259):

Yes?